OpenStudy (anonymous):

Can you help me solve this integral: sin^2t*cos^2t dt

OpenStudy (amistre64):

you can rewrite it if you want using trig identities, might make it simpler

OpenStudy (amistre64):

\[s^2(t)c^2(t)=(s(t)c(t))^2=s^2(\frac12t)\] if i recall that correctly

OpenStudy (amistre64):

err, sin(t)cos(t) = 1/2 sin(2t) soo close

OpenStudy (anonymous):

are you sure? i dont understand about last par

OpenStudy (amistre64):

if you dont understand that identity then maybe: cos^2=1-sin^2 therefore sin^2 cos^2 = sin^2(1-sin^2) = sin^2 - sin^4

OpenStudy (amistre64):

the integrals can be worked out by hand using the by parts method, or by table ....

OpenStudy (anonymous):

so we have integral sin^2 2t dt, right? and 1/4 before integral

OpenStudy (amistre64):

yes, if thats the route you want to take :)

OpenStudy (raden):

that's right, but before or after is same :)

OpenStudy (amistre64):

\[\frac14\int sin^2(u)~du\]if you want to clean it up some

OpenStudy (amistre64):

u = 2t du = 2 dt dt = du/2 sooo \[\frac18\int sin^2(u)du\]

OpenStudy (amistre64):

gotta run, good luck

OpenStudy (anonymous):

but what to do with sin^2? i dont know..

OpenStudy (amistre64):

sin^2 is a basic integral by parts, which is also tabled up

OpenStudy (anonymous):

oo, ok, thank you :)

OpenStudy (raden):

or u can use the identity : sin^2 (2t) = (1 - cos(4t))/2

OpenStudy (anonymous):

thank you too :)